OpenStudy (anonymous):

calculate the average rate of change for the function f(x)=-x^4+4x^3-2x^2+x+1 from x=0 to x=1

OpenStudy (agent0smith):

Why did you never try showing any working in the last question??

OpenStudy (agent0smith):

We aren't just going to give you the answer if you put in no effort.

OpenStudy (anonymous):

i am putting in effort and i dont really appericiate you saying other wise

OpenStudy (agent0smith):

\[\Large f(1)=-(1^4)+4(1)^3-2(1)^2+1+1 = ???\]then work this out.

OpenStudy (anonymous):

13

OpenStudy (agent0smith):

Show every single step of what you're doing.

OpenStudy (anonymous):

i just add and subtract when im they say too

OpenStudy (agent0smith):

Show me what you are adding and subtracting.

OpenStudy (anonymous):

the numbers you posted for me too add and subtract

OpenStudy (agent0smith):

Then show me how you got 13. Show your work. Don't just say the answer is 13, show your steps.

OpenStudy (agent0smith):

First, what is \[\large -(1^4)\]equal to?

OpenStudy (agent0smith):

@helpineedit...???

OpenStudy (anonymous):

sorry never mind i called my teacher and they helped me

OpenStudy (agent0smith):

Do you understand what they explained to you?

OpenStudy (agent0smith):

Because either your answer of 13 was a random guess, or you're doing something very strange when calculating that number.

OpenStudy (agent0smith):

Hence why I asked you to show how you got 13...

OpenStudy (anonymous):

lol yea i was doing it way wrong but ik now so yea

OpenStudy (agent0smith):

That is why I kept asking you to show your work. You never showed ANY work in this entire question or the other one.

OpenStudy (agent0smith):

"Show every single step of what you're doing." helpineedit "i just add and subtract when im they say too" "Show me what you are adding and subtracting." helpineedit "the numbers you posted for me too add and subtract" Do you not see an issue here? show your work means show what you are doing. Not say what you are doing.

OpenStudy (agent0smith):

For example: \[\Large f(1)=-(1^4)+4(1)^3-2(1)^2+1+1 = \] \[\Large -1+4(1)-2(1)+1+1 = \] \[\Large -1+4-2+1+1=\]